1st half of night: NON-sidereal SW testing: 10 January 2012 UT

Observer: CWoodward, KBoutsia (Tucson)
Telescope Operator: DGonzalez (LBTO)
Software Support: MDelapena, CBiddick, NCushing (Tucson)
Telescope Support: (LBTO)
Instrument Support: JMorris (LBTO)

2nd half of night:

Observer: Arakich, OKuhn, KBoutsia (Tucson)
Telescope Operator: DGonzalez (LBTO)
Telescope Support: (LBTO)
Instrument Support: (LBTO)

Summary

Problems with the air pressure on the SX side at the beginning of the night. 0.5hr lost.(?)

1st half we tried SW for NS guiding with LBC. There have been some problems with the guiding in stare mode which were fixed and we still did not manage to run a dither pattern with correct guiding. No time to switch to MODS for testing NS.
We actually had high winds (close to the limit) for part of the NS testing with LBC. The seeing was ranging between 1.5'' and 2.0''

2nd half went well, apart from significant time lost (1.5hr) to some software problem with guiding at the start of proceedings. Corastig sems to work consistently and accurately over a wide range of magnitudes (tested 10th-16th). Next to decide if and how to implement it.

Details

00:44 10Bino_Bias

01:05 lbckill/lbcstart to make sure that the correct config files are loaded.

01:06 Problems with the SX air pressure - call Michael

01:43 Problem seems to have been solved and we are about to start

01:57 Preset to WT10_242 for pointing correction, collimation and rangebal
  • dimm seieng 2.0'' going donw to 1.5''
  • dofpia, /First
-15651      0      0      0      0      0      0   1.10   6.17     red
-18104      0      0      0      0      0      0   1.74   7.13    blue

-13771      0      0      0      0      0      0   1.26   5.43     red
 -1144   -153    123    446    197    -55    -30   1.49   0.51    blue

 -1302    -20    200  -1511    683    -45     14   1.44   0.91     red
 -657   -169    -47    421     85    112    -11   1.59   0.34    blue

  1757    566   -420   1957   1153     73     11   2.46   1.25     red
 338    205   -206    239     -4   -126    -17   1.54   0.22    blue

 -1284   -225    305  -1477   -387   -416    -76   1.19   0.90     red
-956   -236   -176   -271    -71     37    -16   1.45   0.41    blue
  • dofpia, /First, /X2
1578    130    217   1582    271   -520   -100   1.16   1.01     red
2352   -200      9   -338   -389   -483    -96   1.31   1.01    blue

-248   -114    229    531     90    -94    -73   1.28   0.29     red
  595    -14    207    169    217   -275    -71   1.32   0.33    blue

-244    -36     49   -663   -107   -144    -43   1.07   0.34     red
-1096    -50    170     67    115   -149    -26   1.38   0.46    blue

 896    -83    -63   -560    242   -133    -21   1.31   0.46     red
 195    -25   -101     96     22   -135    -34   1.46   0.15    blue
  • lbcrangebal

02:18 Set NSIGUI - send 2006T1Levy_dither.ob
  • NSIGUI name of the ephemeris file is : P2006T1_2012-Jan-11.txt
  • it had to extrapolate value for the ephemeris because the Starting Date/Time is 2012-Jan-11
  • we got a warnig!!! - we should create the NSIGUI coords file again
  • images actually look good!!! and we keep the same file for now
  • 2 things we learned - the OB can have dummy coords since the nsigui overrides them and sends the ephemeris coords - it also extrapolates correctly if you create the ephemeris for the wrong date!!
  • the second lbc red image had a jump!
  • something weird is happening with the guiding - it seems to be jumping around and there are funny trails
  • images with the most obvious problems: lbcr.20120110.022757.fits / lbcb.20120110.022816.fits - we abort

02:33 updated ephemeris to the correct date so no extrapolation is needed
  • Set NSIGUI - P2006T1_2012-Jan-10.txt
  • 2006T1Levy_stare.ob - to see if we still get the jumpy trails

02:40 Abort OB because we still get a jumpy trail and we need to figure out why
  • in the meantime the TO needs to measure the wind.
  • LBC is not sending guiding updates - actually there was no star on the guide-tech-chip

02:49 Slew to 2006TLevy - dofpia
  • we actually need to change the coords in the stare.ob because we need to run dofpia before and we would like to do so close to the target!
  • dimm seeing 1.5''
 3989     23    112    668   -456     32     -8   1.23   1.62     red
 10    -96    105    206   -345      5    -13   1.34   0.20    blue

 1283     32     66    102     97   -195    -43   0.97   0.53     red
  64   -112    -53   -180   -150   -107     14   1.36   0.15    blue

 -1195    -75   -100    -38    -81     53     19   1.30   0.48     red
  434      0   -105      2    150    103     16   1.52   0.21    blue

 -829    -70     45   -101    185    209     79   2.06   0.37     red
 -914     11   -216   -163    166    251     70   2.06   0.42    blue
  • exited and start dofpia with 16sec exp time
 938     92   -114    -36    -71   -195    -36   1.13   0.40     red
 1327    112     48    242   -280   -238    -58   1.62   0.58    blue

-1894    117     62  -1196   -276    179      9   1.31   0.94     red
 728    -96     27     -4    139    -67    -14   1.64   0.31    blue

 1171   -122     35   -379   -387    154     39   1.49   0.54     red
 -2199    -29     87    383    323    335     90   2.14   0.92    blue

  -51    476    107    706    560    118     50   1.49   0.44     red
 -968   -144   -119   -273   -131    202     22   1.47   0.44    blue

 -28    -68     37   -272    253   -144     -6   1.19   0.21     red
-90     47     68    -62   -349    -23    -10   1.16   0.18    blue

03:08 Send the LBC stare ob again WITHOUT guiding
  • set NSIGUI - TURN OFF guiding on LBC (this time) - start the 2006T1Levy_stare.ob
  • the trails still seem a bit jumpy and now we are in open loop tracking
  • the comet does not have fragments following so the previous images with the two tails had guiding issues!
  • in the last images on both sides the trails looked nice and smooth again

03:23 Send the LBC stare ob again WITH guiding
  • set NSIGUI - TURN ON guiding on LBC (this time) - start the 2006T1Levy_stare.ob
  • now with the guiding ON we have the same problem again - it jumps and the comet has double tail and it gets smeared out!
  • abort OB and try to figure out how to fix this.

03:42 send the OB again with updated coords for dofpia
 3722    179    -10   1315   -491     52      2   1.22   1.60     red
 1897    -75     46     53    257   -152    -31   1.41   0.77    blue

  472   -130    -74   -133   -101    -28     16   1.38   0.22     red
 152   -110    -28     60    -41    -34      3   1.40   0.11    blue

 254    -91   -125   -626    495    156     19   1.35   0.40     red
 303    196    -34     96    237    -64    -21   1.34   0.20    blue

03:53 stop and start PCS after implementing some changes trying to correct the guiding problem

03:55 set NSIGUI - guiding is ON - send 2006T1Levy_stare.ob
  • same kind of problem again
  • it seems to be making the jump before the guiding corrections actually start

04:11 SW group is trying another solution - in the meantime we run dofpia
-667   -102    161    944     34     26      4   1.47   0.51     red
 327     82     80   -120     43     -3     11   1.85   0.16    blue

 -698      1    109    -93   -314    212     52   2.15   0.35     red
 -600   -184     91    -70    135    209     31   2.13   0.30    blue

 1158     28   -257    227   -281   -187    -26   1.43   0.51     red
  • dofpia again since tcs was restarted
 975    -86    124   -521    131    111     22   1.45   0.47     red
 393    -17   -135    -42   -252     84     -9   1.48   0.22    blue

 -602   -184    -36   -244     -8    116     32   1.46   0.29     red
  -649   -166     21    -64     72    195     51   1.66   0.30    blue

  984    -91     94    345   -249     87    -10   1.43   0.44     red
 -337    -47    -84    -45     54    145     11   1.60   0.18    blue

-208    104     38   -457    837     97     27   1.58   0.45     red
   79   -124    -95     99    -97     74     19   1.65   0.12    blue

 -845    -98    -71    812   -138    177      7   1.27   0.52     red
 201     65     80    -24     38    -31     -4   1.45   0.12    blue
  • it is not converging on the red and we abort for now...

04:27 set NSIGUI - guiding is ON - send 2006T1Levy_stare.ob
  • this time comment out first guide push so it is not applied
  • now the trails look smoother
  • it seems as if the zero guide correction is causing the jump
  • abort this OB since it seems to be working

04:37 set NSIGUI - guiding is ON - send 2006T1Levy_dither.ob
  • to see if it also works with the dither pattern
  • here it seems to be working on some images and it gives jumpy trails on some others..
  • SW does not understand why...

04:59 Slew to Daphne41 - dofpia
 6710      0      0      0      0      0      0   1.06   2.65     red
 6524      0      0      0      0      0      0   1.28   2.57    blue

 331    102   -151   -316   -982    -53    -36   1.00   0.49     red
 -139    192     10   -291   -149     38     14   1.27   0.19    blue

 -165    157    -80   -105   -118     30     -9   1.12   0.14     red
  250    -23   -150   -252    -81     62     26   1.34   0.18    blue
  • the rates on this asteroid RA=-33''/hr and Dec=+6.5''/hr which is a factor of 10 less than the previous comet
  • check the wind before running the OB

05:11 set NSIGUI - guiding is ON - send 41Daphne_stare
  • we send the telescope to Dec = -00:32:30 but the telescope went to Dec=00:32:30
  • the blue image seems fine - no trails since the target is moving slowly...
  • the red image is completely saturated - are we pointing to the moon???

05:18 set NSIGUI - guiding is ON - send 41Daphne_dither
  • abort because this target is too close to the moon and the images are saturated
  • dimm seeing 1.1.''

05:29 Slew to new target - Victoria_dither.ob
  • no collimation here because we are more interested to see if guiding is behaving correctly
  • this is even slower than previous target

05:30 set NSIGUI - guiding is ON - send Victoria_dither
  • with the slow moving objects the dither seems to be working

05:51 We are changing target and this time we manually set the rates in the NSIGUI
  • we use the differential rates tab and we need to fill in ALL the boxes (including Time) with the correct values.
  • set NSIGUI - send AH06D16_stare.ob
  • aborth OB again because both chips are saturated - decrease exptime

06:00 repeat OB with 90sec exptime
  • we just need to press set in the differential rates tab because once we have a time associated with the coords it extrapolates

06:08 We do not know what the target is because it is very faint
  • we slew to a pointing star to see what are the exact pixel coords where the star shows up to see where the target would go...
  • dofpia
 3379     76   -102    334    153    146     74   1.75   1.35     red
  2060    145     -7    192   -104    201     96   1.89   0.83    blue

  891     68   -132   -101    -61    -24     40   1.09   0.37     red
  1208     95     13     -8    -29    -23     44   1.36   0.48    blue

  499     46    -67    261     32     12     56   1.35   0.24     red
  -480      6     84     -5     15     70     69   1.60   0.21    blue

 70    -14   -100   -136     31     -6     54   1.41   0.11     red
 -471   -100     33     94     79     29     59   1.57   0.21    blue

06:19 set nsigui differential rates - AH06D16_stare.ob
  • we still don't see the asteroid

06:28 to test high rates we put in the differential tab: RArate=-150arcse/hr and Decrate=+200arcsec/hr to see if we get the jumpy guiding
  • we still see the 'peanut' trail but they could be actual guiding corrections.

06:37 set nsigui differential rates - AH06D16_dither.ob
  • we still have the same guiding problem
  • we repeat OB lowering the rates

06:46 set nsigui differential rates - AH06D16_dither.ob
  • now the rates are: RArate=-113arcse/hr and Decrate=+150arcsec/hr
  • the trails look smoother here...
  • the red is not guiding and the trail is smooth - in the blue we still see some dots since it keeps sending guiding corrections

07:02 Send the Daphne OB with the negative Dec to see if it handles it correctly now that the fix has been implemented
  • it did not come through - something is wrong
  • the second time it DID come through with the correct Dec
  • it works!!!

Hand over to AR and reconfiguring to MODS (07:05)

07:20 waking up MODS
  • mods1 start gui
  • execMODS ~MODSeng/modsScripts/Support/modsWake.pro

07:45 testing corastig on MODS
  • on-axis BS star to test basic functionality.
  • first functionality test works! but we cannot do it with track presets because these retract probe. Must send instead an acquire preset and DGH pressed "Stop Guiding" on GCS gui.
08:13
  • First test: collimated, then stopped WFS, sent +10000 nm Z6, corastig corrected with -9684 nm. Some problem with Olga's script introducing focus but backing out too much.
08:17
  • RA 7:34 Dec 29:51:11.9, guiding failing, are we pointing too near the large moon? Checking wind, sounding strong enough to worry Dave. Presetting to WT10_273, GS 1.
08:20
  • Olga thinks she's found the focus back out problem, will test...
  • Still guiding problem, watching guide star rotate about hotspot, calling Tom Sargent to get him to roll back rotator changes as first attempt, then we'll resend preset to test... if that fails will call Michelle to roll back PCS changes. 09:40
  • here's a summary of troubleshooting AGw guiding on MODS.
  • 1) Got Tom Sargent to roll back his rotator software to previously working version. No joy.
  • 2) Got Michelle to roll back PCS to previously working version, guiding still not working. No joy.
  • 3) Michelle also verified that GCS was set back to the same version by Torsten, she copied it over again, though, aqnd the problem seemed to go away. We get back on sky at 9:54.
09:54
  • Adjust pointing, now presetting to GS 7 on ACT0237. First ACTIVE then after converging, resend as ACQUIRE, then "STOP GUIDING" on GCS GUI (which stops the acquisition image movie) and run corastig again with +10000 nm Z6.
  • Corrected astigmatism by -8500 nm.
  • Now trying exceeding the ellipticity limit on the first image, sending +20000 Z6, didn't work so well in this extreme case, only backed out 5000 nm Z6 after many iterations of increasing focus offset to reduce total ellipticity.
  • Tryed it again with ellipticity limit reset from 2.8 to 5. This time it worked without too many focus offsets, overshot the target by 50%, then self corrected on running again. 10:35
  • Now with the higher ellipticity limit in, trying the +10000 nm Z6 initial value. Correction from corastig -24000 Z6 ( frown, sad smile ). Second iteration undid the overshoot. Seems the overshoot is caused by Z6 step of 4000 nm not being sent (essential to corastig algorithm).
  • Now sending -10000 nm Z6 and running corastig. This time corrected +9552 nm in one step.
  • Repeating the +10000 nm Z6 test to see if something systematic is going on, Olga is puzzled by extra guide image. Seeing 2.2 arc seconds. Correcte -13000 nm Z6.
10:55
  • Now doing a test where we alternately preset alternately to high and low stars. First we will collimate up high, then preset 5 times, low, high low etc. running corastig then sending an active preset at each preset. Once converged at each star on to the next preset. We time the total time it takes to run corastig then converge on 5 stars. Then repeat the test without corastig.

11:07 Presetting to WT10_350. Cleaning pointing then sending active preset. Using WT10_350 is our high el star, GS 1 (mag 13.6). WT10_411 is our low el star, GS 2 (mag 12.2).
11:19 Start test, preset acquire to WT10_411, GS 2. Run corastig, run active optics, converged.
11:29 preset to WT10_350, acquire, stop guiding, run corastig, guide probe didn't move. Time out from 10:31 - 10:32. Running corastig. Sending active preset.
11:36 preset to WT10_411 GS2 acquire mode. Run corastig. Running active preset.
11:43 preset to WT10_350 GS1, acquire, run corastig. run active preset to same star.
* test finished. 11:48.
11:49 New non-corastig test. active preset to low star.
11:57 active preset to high star.
12:03 active preset low star.
12:10 active preset to high star.
12:17 finished.
  • redoing corastig test.
12:18 preset to low el star. run corastig, run active preset.
12:25 preset to hi el star, run corastig, run active preset.
12:32 preset to low el star, run corastig, run active preset. Guiding bad. First WFS image gave 800 nm, then 500, then 500, then 375 etc.
12:38 preset to hi el star, run corastig, run active preset.
12:44 finished.
12:50 Now running corastig on GS 2, WT10_350, a 14.95 mag star, Corastig worked well.
12:55 now trying a mag 16 star GS 6 on WT10_350. First collimating down low.
01:17 That worked well, now trying a 11th magnitude star, GS 4, first collimating down low.
01:24 Note that there are two stars close together here, might spoil things. Yes it did, the two stars were orientated such that they added to make the ellipse look like a circle.
01:31 Trying again oin mag 10 star @ ACT0339 (GS3)
01:36 Worked, so corastig is giving similar results over magnitude range of 10-16. Nice. Our work here is done.

-- KonstantinaBoutsia - 10 Jan 2012
Topic revision: r9 - 11 Jan 2012, KonstantinaBoutsia
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback